OpenStudy (anonymous):

I'LL GIVE MEDAL!!!! What value of z makes the equation true? 1/6 = z-7/12

OpenStudy (trojanpoem):

Add 7/12 to both sides z = 2/12 + 7/12 = 9/12

OpenStudy (anonymous):

that doesn't really make sense... where do you get 2/12?

OpenStudy (anonymous):

go here i hope it helps you......im really not that good at fractions.... sorry. https://answers.yahoo.com/question/index?qid=20081201092945AAGxfM4#

OpenStudy (anonymous):

ya...no... their is a question close to that, but not with same numbers... thanks though!

OpenStudy (trojanpoem):

@StarClia , You don't make sense 1/6 = 2/12 ( Algebra 1 , for metal summation)

OpenStudy (anonymous):

okay.....im sorry buddy..... hold on

OpenStudy (anonymous):

oh.. duh... sorry I wasn't thinking...

OpenStudy (trojanpoem):

and 9/12 = 3/4

OpenStudy (anonymous):

okay......im trying to figure it out to......

OpenStudy (trojanpoem):

@kittkatt22 , what are you trying to figure ?

OpenStudy (anonymous):

the equation i need to learn fractions better

OpenStudy (anonymous):

what z is.... so... z is 9/12? but that doesn't work... oh! I think I know... Do I divide 12?

OpenStudy (trojanpoem):

1/6 = z-7/12 z = 1/6 + 7/12 -> add this to the calc or make a common denominator.

OpenStudy (trojanpoem):

z = 9/12 9 = 3* 3 12 = 3 * 4